OpenStudy (anonymous):

if 200 cm3 of a gas of 0.268 g at STP . a) calculate its molar mass b) calculate its volume at 34 degree Celsius and 950 torr

OpenStudy (jfraser):

What you have is a known volume, temperature, and pressure of a gas. The ideal gas law says:\[P*V = n*R*T\] where R is constant (0.08205L*atm/mol*K) Rearrange the equation to solve for moles:\[n = \frac{P*V}{R*T}\] make sure your volume is in Liters, plug in your values, and find moles. Molar mass is:\[MM = \frac {g}{mol}\] the problem gives a mass, and you've just calculated moles. Divide & solve.

OpenStudy (anonymous):

i still don't understand . how come I have just calculated the moles ? . and 200 cm3 is ?

OpenStudy (jfraser):

molar mass has units of grams per mole. You need to know how many moles of gas you have, in order to find the molar mass. 200cm^3 is the same volume as 200mL, which you can convert into liters

OpenStudy (anonymous):

okay thanks . such a help

OpenStudy (anonymous):

can i ask you another question ?

OpenStudy (jfraser):

sure

OpenStudy (anonymous):

okay . here it is if 5600 cm3 of a gas has a mass of 11.0 g at STP , what is its relative molacular mass ?

OpenStudy (jfraser):

same question. you have a mass, a volume, and STP. Plug in to solve for moles, then divide mass by moles to get molar mass. THe numbers are different, but the problem is exactly the same

OpenStudy (anonymous):

ooo , act i kinda nervous when i saw rmm . bcoz , i think it is differ than molar mass . it is just the same right ?

OpenStudy (jfraser):

essentially, yes. It's relative to hydrogen, which has a mass of 1 anyway

OpenStudy (anonymous):

ohh okay , i got it now . thank you so much ! :D

OpenStudy (anonymous):

there is one question left a 1.58 g sample of c2h3x3 (g) has a volume of 297 mL at 769 torr and 35 degree Celsius . identify element x

OpenStudy (jfraser):

Do the same thing as before. Find the molar mass of the compound (C2H3X3) using the ideal gas law (PV=nRT) Subtract the mass of 2-carbons. Subtract the mass of 3-hydrogens. Divide the remaining mass by 3. On the periodic table, find the element with a mass equal to that number.